PE 10-1A On October 12, Belleville Co. borrowed cash from Texas Bank by issuing a 30-day note with a face amount of $70,000,ice 11... i need 1a, 2a, 3a solved
Answer #1

Solution 10-1A

Proceeds from an interest-

bearing note are given by

Face value =P(1+r × d/365)

Face value =$70,000 (this amount will be need to pay after maturity)

P = the discounted amount we get upfront =?

d =time to maturity in days =30days

P =(1+0.6×30/365) =$70,000

P=$70,000/1.0493151

P=$66,710

b.) Proceeds from discounted note

=face value of the note - discount

Discount on a short term note

=face value × discount rate ×days to maturity/365

$70,000×6%×30/365 =$345

Proceeds =$70,000-$345

=$69,655